### v2

# IPO: Iterative Preference Optimization for Text-to-Video Generation

Abstract
Video foundation models have achieved significant advancement with the help of network upgrade as well as model scale-up. However, they are still hard to meet requirements of applications due to unsatisfied generation quality. To solve this problem, we propose to align video foundation models with human preferences from the perspective of post-training in this paper. Consequently, we introduce an Iterative Preference Optimization strategy to enhance generated video quality by incorporating human feedback. Specifically, IPO exploits a critic model to justify video generations for pairwise ranking as in Direct Preference Optimization or point-wise scoring as in Kahneman-Tversky Optimization. Given this, IPO optimizes video foundation models with guidance of signals from preference feedback, which helps improve generated video quality in subject consistency, motion smoothness and aesthetic quality, etc. In addition, IPO incorporates the critic model with the multi-modality large language model, which enables it to automatically assign preference labels without need of retraining or relabeling. In this way, IPO can efficiently perform multi-round preference optimization in an iterative manner, without the need of tediously manual labeling. Comprehensive experiments demonstrate that the proposed IPO can effectively improve the video generation quality of a pretrained model and help a model with only 2B parameters surpass the one with 5B parameters. Besides, IPO achieves new state-of-the-art performance on VBench benchmark. We will release our source codes, models as well as dataset to advance future research and applications.
Figure 1: The overview of iterative preference optimization(IPO) framework
1 Introduction
Text-to-video generation has drawn lots of attention due to its great potentials in movie products, video effects and user-generated contents, etc. Recent works [3, 20, 13] introduce the Transformer architecture into diffusion models and improve the video generation quality by simply scaling up the parameter size, e.g., CogVideoX-5B [52], Mochi-10B [44] and HunyuanVideo-13B [19]. Although achieving impressive performance, they still face challenges for generating user-satisfied videos with consistent subject, smooth motion and high-aesthetic picture. In addition, their large model sizes make them slow to produce videos. These drawbacks limit their applications in practice.
To align generation results with human preference, post-training techniques have shown their effectiveness in fields of Large Language Models (LLMs) and Text-to-Image models (T2Is). For instance, Direct Preference Optimization (DPO) [36] exploits pairwisely ranked data to make LLMs know the language style liked by human. While Kahneman-Tversky Optimization (KTO) [8] uses pointwisely binary data to reinforce T2Is to learn how to maximumly align the objective with human expectation. These techniques promote generative models to derive user-satisfied contents and give us inspiration for improving video generation models.
Thus, we propose to align Text-to-Video models (T2Vs) with human preference from the perspective of post-training in this paper. In particular, we present the Iterative Preference Optimization (IPO) framework for enhancing T2Vs. IPO introduces preference priors for finetuning T2Vs in the form of reinforcement learning, which considers subject consistency, motion smoothness and aesthetic quality, etc. This leads to improved video generations that are well aligned with human preference. In addition, difference from prior single-round methods, IPO adopts a multi-round optimization strategy to reinforce T2Vs in an iterative manner, which strengths the models continuously. In this way, IPO can effectively improve T2Vs without need of large-scale dataset and expensive supervised finetuning.
Specifically, IPO consists of three core parts as shown Fig. 1: the preference dataset, the critic model and the iterative learning framework. IPO first collects a preference dataset for training the critic model. To get this dataset, IPO predefines several categories, including human, animal, action, etc. Then, IPO randomly combines these categories and uses LLM to extend them as prompts for generating videos with T2V models. Given these videos, IPO manually annotates them with two kinds of labels: one is pairwise ranking label that depicts preference comparison of two videos; another is pointwise scoring label that describes video quality as good or bad in a binary manner. After collecting the preference dataset, IPO uses it to train a critic model for justifying the quality of input videos. Concretely, IPO exploits Multi-modality Large Language Models (MLLMs) to achieve the above goal considering their powerful capability for recognizing video contents. In addition, IPO combines instruction-tuning to learn the critic model, making it able to generate pairwise and pointwise labels with different prompts. Given the critic model, IPO plays as a plugin for preference optimization of T2Vs without requirements of retraining or relabeling. For a specific T2V model, IPO can either use Diffusion-DPO [45] with pairwise ranking data or Diffusion-KTO [22] with pointwise scoring data. In addition, IPO adopts the critic model for automatically labeling videos from the reinforced T2Vs, which can be used as prior information for further improving T2Vs. In the iterative way, IPO can consistently improve T2Vs for generating high-quality videos.
Comprehensive experiments on benchmark VBench demonstrate that the proposed IPO can effectively and efficiently reinforce T2Vs to produce user-satisfied videos with consistent subject, smooth motion as well as high-aesthetic quality. In addition, given CogVideoX as the video foundation model, IPO can improve a 2B model to surpass the 5B counterpart. Our contributions can be summarise into three folds: First, we introduce a new preference optimization framework into text-to-video generation, which well align generations with human preference. Second, we present a novel iterative post-training framework for enhancing foundation models in a multi-round manner, differing from prior single-round methods. Third, we set new state-of-the-art on the benchmark VBench both in quantity and quality. To advance future researches in the area of video generation, we will release the preference dataset, the codes, the critic models as well as T2V models.
2 Related Work
Text-to-Video Generation
Recent advances in video generation have been predominantly driven by diffusion models [16, 40, 41, 55], which have demonstrated remarkable improvements across multiple dimensions, including diversity, fidelity, and image quality [5, 7, 17, 26, 38, 47, 49], with the scaling of training data and model size further enhancing their performance [14, 39, 48]. In the domain of Text-to-Video (T2V) generation, pre-training models from scratch demands substantial computational resources and access to extensive human-curated video datasets, highlighting the critical importance of leveraging large-scale datasets for effective model training [10, 29, 20, 3, 19]. Current methodologies, such as those presented in [2, 46, 12], build upon text-to-image models by incorporating spatial attention mechanisms, introducing temporal attention modules, and fine-tuning them on video datasets, often employing joint image and video training strategies [30], with [12] proposing a plug-and-play temporal module that facilitates video generation from personalized image models. State-of-the-art results in terms of pixel quality and temporal consistency have been achieved by methods such as [56, 4], while recent diffusion models based on the Dit structure, such as [3, 52, 4], have demonstrated exceptional performance, advancing the frontiers of T2V generation by enhancing spatiotemporal coherence and scalability, showcasing the potential of Transformer-based diffusion architectures. Further advancements have been made in addressing the discrepancy between spatial and temporal modules in two-stage training, with works such as [24, 19, 52] employing 3D full attention mechanisms to improve overall performance and effectively bridge the gap between spatial and temporal components.
The training of large models typically involves two key stages: pre-training and post-training, with relatively limited research focusing on the post-training phase for text-to-video (T2V) generation. In this context, [53] introduced a temporal decaying reward (TAR) mechanism, grounded in the assumption that the central frame of a video should be assigned the highest importance, while the emphasis gradually diminishes towards the peripheral frames. This strategic allocation of importance across frames ensures a more stable and visually coherent video generation process. Meanwhile, [21] explored the extensive design space of energy functions to enhance ordinary differential equation (ODE) solvers, demonstrating its potential by improving T2V model training through the extraction of motion priors from training videos. Additionally, VADER [35] proposed a method that leverages the gradient of a reward model to fine-tune a base video diffusion model, utilizing a wide range of pre-trained vision models to align various video diffusion models. This alignment approach, facilitated by VADER, exhibits strong generalization capabilities, even for cues not encountered during training, thereby advancing the adaptability and performance of T2V systems.
Aligning with human preferences
Aligning human preferences through post-training has been widely studied for large language models (LLMs), with several approaches [31, 37, 59] making notable advancements. Reinforcement Learning from Human Feedback (RLHF) [31] trains a reward function from comparison data on model outputs to align the policy model. Rank Responses for Human Feedback (RRHF) [54], a simpler alternative to Proximal Policy Optimization (PPO), extends supervised fine-tuning and reward model training. Sequence Likelihood Calibration (SLiC) [58] achieves alignment using human feedback data collected for different models, akin to offline RL data. Direct Preference Optimization (DPO) [36] simplifies alignment by directly training on human preference data, improving stability and performance while reducing computational costs. Kahneman-Tversky Optimization (KTO) [8] maximizes utility based on the Kahneman-Tversky model, rather than maximizing log-likelihood of preferences. Rejection Sampling Optimization (RSO) [28] introduces statistical rejection sampling to derive preference data from the optimal policy, enhancing alignment with human preferences.
Several methods have successfully integrated preference alignment with diffusion models, driving significant progress in the field. DraFT [6] fine-tunes models to maximize differentiable reward functions, such as human preference scores, while [34] aligns models with downstream reward functions via end-to-end backpropagation through the denoising process. Similarly, [50] optimizes models based on reward scores, but these approaches often face inefficiency and instability. To overcome these challenges, DPOK [9] frames fine-tuning as a reinforcement learning problem, combining policy optimization with KL regularization and policy gradients to update pre-trained text-to-image diffusion models. DDPO [1] reframes denoising as a multi-step decision problem, using policy gradients to enhance cue-image alignment without additional data collection or human annotation. Diffusion-DPO [45] optimizes models using human comparison data, reformulating Direct Preference Optimization (DPO) to improve model likelihood and differentiable objectives. Diffusion-KTO [22] introduces a novel approach for aligning text-to-image models by maximizing expected human utility, removing the need for costly pairwise preference data. The iterative preference optimization algorithm in this work further enhances offline learning methods like Diffusion-DPO and Diffusion-KTO, addressing instability and performance issues while advancing alignment effectiveness.
3 METHOD
In this section, we present our proposed general iterative preference optimization (IPO) framework, which is designed to support both pairwise and pointwise preference alignment optimization algorithms. For offline preference alignment algorithms, the effectiveness is often hindered by the limitations of offline datasets, as highlighted by RSO [27], particularly when the offline data distribution diverges from the target distribution. Taking Direct Preference Optimization (DPO) [36] as a representative example, which can be extended to other algorithms, DPO can be interpreted as imposing constraints on the resulting policy through the dataset $D$. Informally, for DPO to converge to the optimal policy $\pi^{*}$, it requires an infinite dataset $D$ to comprehensively cover the entire query-response space. However, in practice, training on a finite dataset $D$ inevitably fails to encompass the full query-response space, leading to suboptimal convergence and performance limitations, which our IPO framework aims to address by iteratively refining the alignment process and enhancing the coverage and quality of the policy optimization.
[23] demonstrates that reward models may exhibit superior generalization capabilities compared to policies in terms of sample complexity, meaning that, given a fixed number of samples, reward models achieve higher preference classification accuracy. To leverage this insight, we first train a critic model capable of performing both pairwise ranking and pointwise scoring. During the training process, no additional human feedback is queried; instead, the trained critic model is utilized to label the samples generated by the model in subsequent training iterations. However, since the model is initially poorly aligned with human feedback, the generated results often suffer from low quality, leading to inefficient sampling. This results in collected samples that may not adequately approximate the target distribution, ultimately causing performance degradation when training on these samples. To address this issue and enhance algorithmic efficiency, we propose an iterative preference optimization (IPO) framework, which progressively refines the model’s alignment with human preferences, enabling a more effective and gradual achievement of the final optimization goal.
In RLHF, the optimization goal is to maximize the expected reward given the condition $c$, while increasing the KL-divergence as a reward penalty to keep the optimized model close to the reference distribution.
$\max_{\pi_{\theta}}\mathbb{E}_{x_{0}\sim\pi_{\theta}(x_{0}|c)}\left[r(x_{0},c)-\beta\log\frac{\pi_{\theta}(x_{0}|c)}{\pi_{ref}(x_{0}|c)}\right],$ (1)
where the hyperparameter $\beta$ controls regularization and is usually a fixed value.
In [57], a loss functional is defined in terms of $\pi_{\theta}(\cdot|c)$, expressed as
$\mathbb{E}_{x_{0}\sim\pi(x_{0}|c)}\Big{[}-r(x_{0},c)-\beta\log\frac{\pi_{ref}(x_{0}|c)}{\pi_{\theta}(x_{0}|c)}\Big{]}=\\ \beta\mathbb{KL}\Big{(}\pi_{\theta}(x_{0}|c)\Big{\|}\pi_{ref}(x_{0}|c)\exp\Big{(}\frac{1}{\beta}r(x_{0},c)\Big{)}\Big{)},$ (2)
the value that minimizes the loss functional is $\pi^{*}(x_{0}|c)\propto\pi_{0}(x_{0}|c)\exp\Big{(}\frac{1}{\beta}r(x_{0},c)\Big{)}$. This is consistent with the results of DPO, which directly transforms the optimization objective into training $\pi^{*}(x_{0}|c)$.
The conclusion drawn is that the reinforcement learning process, when combined with the maximum expected reward under KL regularization, aims to approximate $\pi^{*}(x_{0}|c)$ during the learning process. However, in practical scenarios, obtaining unlimited offline datasets to achieve this goal is infeasible, and the model often achieves only limited improvement and constrained generalization when trained on finite and suboptimal offline datasets. While it is necessary to utilize the Critic model to annotate the results generated by the model, the efficiency and effectiveness of data generation often lead to suboptimal outcomes in a single training iteration, highlighting the need for iterative refinement to enhance performance and generalization capabilities.
Inspired by [31], we introduce an iterative preference optimization method that, rather than directly approximating $\pi_{\theta}$ by $\pi_{\text{ref}}\exp\left(\frac{1}{\beta}r(x_{0},c)\right)$—a challenging task—divides the process into multiple stages by examining a series of intermediate distributions. By approximating $\pi_{\text{ref}}\exp\left(\frac{1}{\beta_{i-1}}r\right)$ with $\pi_{\text{ref}}\exp\left(\frac{1}{\beta_{i-1}}r\right)$ at each iteration, we achieve stepwise improvements, making the optimization process more tractable and effective.
$\pi_{ref}^{(0)}=\pi_{ref},$
$\pi_{ref}^{(1)}=\pi_{ref}^{(0)}\exp\Big{(}\frac{1}{\beta_{1}}r\Big{)},$
$\pi_{ref}^{(2)}=\pi_{ref}^{(1)}\exp\Big{(}\frac{1}{\beta_{2}}r\Big{)},$
$\vdots$
$\pi^{*}=\pi_{ref}^{(N)}\exp\Big{(}\frac{1}{\beta_{2}}r\Big{)}.$
To validate the effectiveness and generalizability of the proposed approach, we introduce two distinct offline preference alignment methodologies: Diffusion-DPO, which leverages pairwise ranking data to align models with human preferences, and Diffusion-KTO, which utilizes pointwise scoring data to optimize alignment objectives, both designed to enhance the performance and adaptability of the framework across diverse scenarios.
Diffusin-DPO Objective
Diffusion-DPO [45] extends DPO [36] by introducing inverse decomposition to replace $\pi_{\theta}$ and $\pi_{ref}$. The forward process $q(x_{1:T}|x_{0})$ serves as an approximation for the reverse process $p_{\theta}(x_{1:T}|x_{0})$. Ultimately, the objective function is formulated after these adjustments. By redefining data likelihood in the context of diffusion models, Diffusion-DPO introduces a straightforward and effective loss formulation, enabling robust and efficient preference optimization.
$L_{diffusion-dpo}(\theta)=\\ -\mathbb{E}_{(x^{w}_{0},x^{l}_{0})\sim D,t\sim\mathcal{U}(0,T),x^{w}_{t}\sim q(x^{w}_{t}|x^{w}_{0}),x^{l}_{t}\sim q(x^{l}_{t}|x^{l}_{0})}\\ \log\sigma\left(-\beta T\omega(\lambda_{t})\left(\right.\right.\\ \|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\|^{2}_{2}-\|\epsilon^{w}-\epsilon_{\text{ref}}(x_{t}^{w},t)\|^{2}_{2}\\ \left.\left.-\left(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|^{2}_{2}-\|\epsilon^{l}-\epsilon_{\text{ref}}(x_{t}^{l},t)\|^{2}_{2}\right)\right)\right)$ (3)
where $x^{w}$ is the chosen sample, $x^{l}$ is the rejected sample, $\omega(\lambda_{t})$ a weighting function [15, 18, 36]).
Diffusin-KTO Objective
Inspired by prospect theory, KTO proposes the concept of human-aware loss functions by maximizing the utility function and defining reference points. KTO does not require preference data and can directly use data marked with binary signals to train the algorithm, making it more sensitive to negative samples. Instead of optimizing the expected reward, Diffusion-KTO adopts a nonlinear utility function that calculates the utility of an action based on its value $Q(a,s)$ relative to a reference point $Q_{\text{ref}}$.
$L_{diffusion-kto}(\theta)=\underset{\pi_{\theta}}{\text{max}}\mathbb{E}_{x_{0}\sim\mathcal{D},t\sim\text{Uniform([0,T])}}\\ [U(w(x_{0})(\beta\log\frac{\pi_{\theta}(x_{t-1}|x_{t})}{\pi_{\text{ref}}(x_{t-1}|x_{t})}-Q_{\text{ref}}))]$ (4)
where $w(x_{0})=\pm 1$, $x_{0}$ represents a state that is considered either desirable or undesirable. The quantity $Q_{ref}$ is introduced as $\beta\mathbb{D}_{\text{KL}}[\pi_{\theta}(a|s)||\pi_{\text{ref}}(a|s)]$, which measures the Kullback-Leibler divergence between the policy $\pi_{\theta}$ and the reference policy $\pi_{\text{ref}}$.
Iterative Preference Optimization (IPO) framework, each iteration training consists of three key components: video generation, critic, and preference learning. The implementation details will be introduced in the next section.
4 Implementation details
In this section, we present our iterative preference optimization framework, along with its implementation details and pipeline. An overview of the algorithm is illustrated in Fig. 1. Specifically, the data collection and annotation process are described in detail in Section 4.1, while Section 4.2 elaborates on the training methodology for the Critic model. Finally, Section 4.3 introduces a comprehensive framework for aligning human feedback, encompassing the training of DPO on paired data and KTO on single data.
4.1 Datasets Annotation Pipeline
Video-text collection
Our annotated dataset is constructed using the open-source model CogVideoX-2B [52]. The process begins by creating a diverse set of prompts for video generation, encompassing various character types, actions, scenes (both realistic and abstract), animals, plants, vehicles, and video styles. These elements are randomly combined and refined into coherent prompts using the Qwen2.5 [51] large language model. To ensure diversity and avoid the generation of repetitive sentences, we fine-tune the model’s output by adjusting key parameters such as temperature, top-k, and top-p. Using these prompts, the video generation model produces three distinct variants for each video by applying different random seeds. Throughout the training process, we iteratively regenerate data using an optimized strategy to progressively expand and enhance our preference-aligned dataset.
Human Annotation
When learning to align human feedback, we focus on three dimensions, referring to the text-to-image evaluation work Evalalign[43]: consistency between text and video, video fidelity, and additional motion smoothness. To capture human preferences, we employ two annotation methods: Pairwise ranking and Pointwise scoring.
For pairwise ranking, annotators are presented with two videos generated from the same prompt and instructed to evaluate them based on predefined criteria. During the evaluation, annotators are required to: (1) provide detailed natural language feedback, (2) identify specific quality issues in each video (or confirm their absence), and (3) articulate the reasoning behind their ranking decisions. These annotated comparisons are then systematically organized to construct the training samples:
{<Video1>, <Video2>, Question,Evaluation Dimension,Ranking $\left\{\begin{array}[]{l}\texttt{<Video1>, <Video2>, Question,}\\ \texttt{Evaluation Dimension,}\\ {\color[rgb]{0,0,1}\text{Ranking~{}Answer}},{\color[rgb]{0,0,1}\text{Reason}}\end{array}\right\}$
For pointwise scoring, each video is evaluated across three dimensions, with scores categorized into three levels: good, average, and poor. Similar to the ranking process, annotators are required to provide detailed justifications for their scores, ensuring transparency and consistency in the evaluation process. These annotated comparisons are then systematically organized to construct the training samples:
{<Video>, Question,Evaluation Dimension,Scoring $\left\{\begin{array}[]{l}\texttt{<Video>, Question,}\\ \texttt{Evaluation Dimension,}\\ {\color[rgb]{0,0,1}\text{Scoring~{}Answer}},{\color[rgb]{0,0,1}\text{Reason}}\end{array}\right\}$
Each dataset is annotated by at least three individuals. The final annotation result is determined through voting, selecting the annotation with the highest level of consistency among the reviewers.
4.2 Critic model training
To develop a comprehensive model capable of handling both paired data sorting and individual video scoring tasks, we fine-tuned the pre-trained Vila model [25], leveraging its exceptional capabilities in image and video instruction following, comprehension, and multi-image processing, which make it particularly suitable for training our Critic model. During training, we implemented key modifications, including restructuring the prompt format to incorporate detailed reasoning chains before point-by-point scoring or pairwise ranking, and constructing explicit chains of thought to enhance the model’s decision-making process. This approach enables the model to provide comprehensive rationales prior to delivering its final output, improving both the accuracy of results and the quality of natural language feedback.
We initialized the model using the Vila 13B/40B pre-trained checkpoint and fine-tuned it for 5 epochs on our labeled dataset with a learning rate of 2e-5, distributed across 32 GPUs with a batch size of 4 per GPU. To mitigate inconsistencies from scaling and center cropping, we implemented pad training. For better video content understanding, we extracted 16 uniformly sampled frames from each video sequence, keeping other hyperparameters aligned with Vila’s defaults. This setup optimizes both computational efficiency and model performance, ensuring strong video-text alignment.
4.3 Iterative training pipeline
Iterative preference optimization provides substantial advantages over offline learning methods. The iterative learning framework we propose is a versatile and general-purpose approach, capable of seamlessly integrating various offline learning algorithms into an iterative setting, thereby achieving significant performance enhancements. Specifically, we investigate two distinct offline preference alignment methods: the KTO algorithm, which utilizes pointwise scoring, and the DPO algorithm, which is based on pairwise ranking.
Iterative DPO
In the iterative preference optimization process employing Direct Preference Optimization (DPO), the system follows a structured workflow: during each iteration, the current optimal policy model generates multiple video outputs for a given prompt. These generated videos are subsequently evaluated and ranked by a critic model based on predefined quality metrics. The ranked outputs then serve as training data for preference alignment, where the policy model is fine-tuned to better align with the desired preferences. This cyclic process of generation, evaluation, and optimization continues iteratively, progressively enhancing the model’s performance through successive refinement cycles.
  • Video Generation. We use CogvideoX-2B as the initial checkpoint, and generate at least three random videos for the same prompt by setting different seeds to to form a dataset $\mathcal{V}$ = $\{(x_{k},y_{k_{1}},y_{k_{2}},y_{k_{3}},...)\}_{k=1}^{N}$, where $x_{k}$ represents the prompt and $y_{k}$ represents the corresponding video generated by the model.
  • Pairwise Ranking. To enhance ranking reliability and mitigate potential biases in the Critic model, we implement a robust pairwise comparison protocol with position swapping during inference: for each video pair, we conduct two reciprocal evaluations by alternating their presentation order, retaining only consistent results as valid entries for the preference dataset. Formally, we construct the pairwise dataset as $\mathcal{D}=\{(x_{k},y_{k_{w}},y_{k_{l}})\}_{k=1}^{N}$, where $y_{k_{w}}$ and $y_{k_{l}}$ denote the preferred (winning) and dispreferred (losing) samples, respectively. This rigorous validation process ensures higher data quality and reduces positional bias in the collected preference pairs.
  • Preference learning. During training, we employ a selective sampling strategy to construct high-quality preference pairs by extracting only the highest-ranked and lowest-ranked samples from the critic model’s evaluations, forming the final paired data $(y^{w},y^{l})$ while deliberately excluding intermediate-ranked samples to ensure clearer preference distinctions. For optimization, we utilize the diffusion-DPO loss as the primary objective function and introduce two auxiliary negative log-likelihood (NLL) loss terms following  [33]: the first term, weighted by 0.2, regularizes the preferred samples, while the second term, scaled by 0.1, anchors the model to real video data distributions. This dual regularization strategy preserves the structural integrity and format of generated outputs while preventing undesirable degradation in the log-probability of selected responses, as demonstrated in  [32],  [11], and  [33].
    $\mathcal{L}(\theta)$ $=L_{diffusion-dpo}(\theta)$ (5)
    $\quad+\lambda_{1}\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{sample}}}\bigl{[}-\log p_{\theta}(y^{w}|x)\bigr{]}$
    $\quad+\lambda_{2}\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{real}}}\bigl{[}-\log p_{\theta}(y|x)\bigr{]},$
Iterative KTO
KTO [8] is an offline learning method that leverages pointwise scoring data, eliminating the need for preference-based datasets. By directly utilizing binary-labeled data, KTO effectively trains algorithms while demonstrating heightened sensitivity to negative samples. This makes it particularly adept at handling imbalanced datasets with uneven distributions of positive and negative samples. Furthermore, KTO achieves significant performance improvements even without relying on the SFT stage. In scenarios where data is imbalanced, paired sorting data is unavailable, or preference data is noisy, KTO stands out as a more effective and robust solution.
  • Video Generation. Similar to the DPO training framework, we adopt an identical video sampling strategy in this study. In alignment with the KTO [8] methodology, empirical evidence suggests that sampling multiple videos for a given prompt yields superior results compared to single-video sampling. Therefore, we maintain consistency with the DPO sampling protocol throughout our implementation.
  • Pointwise scoring. The Critic model is employed to evaluate and assign scores to all generated text-video pairs $(y,x)$. Each data sample is categorized into three distinct quality levels: ”Good”, ”Normal”, and ”Bad”. Based on this classification, we designate samples labeled as ”Good” and ”Normal” as positive instances, while those marked as ”Bad” are treated as negative instances for subsequent analysis and model training.
  • Preference learning. We use the Eq. 4 for training. KTO needs to estimate the average reward Q during training. When the batch size is small, the Q value is not accurately estimated and the model training will be unstable. We set the batch size to 128 and achieved good results. To stabilize the training, we also sampled data from the real data VidGen-1M [42] and added negative log-likelihood loss.
    $\mathcal{L}(\theta)$ $=L_{diffusion-kto}(\theta)$ (6)
    $\quad+\lambda\cdot\mathbb{E}_{(\textbf{{y}},x)\sim\mathcal{D}^{\mathrm{real}}}\bigl{[}-\log p_{\theta}(y|x)\bigr{]},$
Both DPO and KTO employ an iterative learning framework that cyclically executes three core steps: (1) strategic data sampling, (2) model-based preference labeling, and (3) alignment optimization. The process incorporates early stopping based on validation metrics while adaptively optimizing strategies for each component to maximize training efficiency and model performance.
5 Experiments
Models
Total
Score.
Quality
Score.
Semantic
Score.
Subject
Consist.
Background
Consist.
Temporal
Flicker.
Motion
Smooth.
Aesthetic
Quality
Dynamic
Degree
Image
Quality
CogVideoX-2B 80.91 82.18 75.83 96.78 96.63 98.89 97.73 60.82 59.86 61.68
KTO 82.47 83.56 78.08 96.85 97.08 99.27 97.93 62.47 68.83 62.31
DPO 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
Table 1: Verify the universality and effectiveness of iterative preference optimization, evaluate the pairwise-based DPO and the KTO algorithm based on pointwise data.
Models
Total
Score.
Quality
Score.
Semantic
Score.
Subject
Consist.
Background
Consist.
Temporal
Flicker.
Motion
Smooth.
Aesthetic
Quality
Dynamic
Degree
Image
Quality
CogVideoX-2B 80.91 82.18 75.83 96.78 96.63 98.89 97.73 60.82 59.86 61.68
IPO-iter1 81.69 82.87 77.01 96.77 97.01 99.01 97.84 61.55 64.33 62
IPO-iter2 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
IPO-iter3 82.74 83.92 78.00 96.79 97.48 99.35 98.17 62.31 69.46 62.87
Table 2: Verify the effect of iterative preference optimization and multiple rounds of training, we evaluated the performance after three rounds of iterative optimization on VBench
Models
Total
Score.
Quality
Score.
Semantic
Score.
Subject
Consist.
Background
Consist.
Temporal
Flicker.
Motion
Smooth.
Aesthetic
Quality
Dynamic
Degree
Image
Quality
CogVideoX-2B 80.91 82.18 75.83 96.78 96.63 98.89 97.73 60.82 59.86 61.68
vila 13b 82.42 83.53 77.97 96.81 97.23 99.21 98.03 62.35 68.76 61.77
vila 40b 82.52 8363 7807 96.83 97.37 99.19 98.09 62.38 68.92 62.01
Table 3: Evaluate the impact of the critic model on performance,verify the performance of the critic model results for different model sizes of 13B and 40B, a more accurate critic model leads to improved results.
Figure 2: Qualitative Comparison. We compare our (iterative preference optimization) performance with CogVideoX-2B.
This section presents the experimental setup, results, and ablation studies to demonstrate the effectiveness of the proposed online reinforcement learning framework for video generation optimization using human feedback during post-training.
5.1 Experimental Setup
Our experiments are conducted on a Critic model constructed based on the diffusion model, which evaluates video quality through semantic and temporal consistency. The Critic model incorporates both metrics derived from human feedback and pre-defined objective criteria to ensure comprehensive evaluation.
The training process leverages both DPO and KTO methods within the proposed online reinforcement learning framework. The DPO method aligns the generated videos with human feedback by minimizing divergence between the learned policy and a reference policy, while KTO directly maximizes the utility of the generated video instead of maximizing the preference log-likelihood. Both methods were implemented with a batch size of 128, learning rate of 2e-5, and trained for up to three rounds of iterative optimization. Each round involves a combination of human feedback-guided rewards and the original diffusion model loss to balance consistency and realism.
5.2 Results
5.2.1 Quantitative Results
Table 4: Comparison of the proposed method with state-of-the-art (SOTA) approaches. Metrics are evaluated using VBench.
Method Quality Score Semantic Score Total Score
Baseline $2B$ 82.18 75.83 80.91
Baseline $5B$ 82.75 77.04 81.61
Optimized $2B$ (Ours) 83.92 78.00 82.74
Models
Subject
Consist.
Background
Consist.
Temporal
Flicker.
Motion
Smooth.
Aesthetic
Quality
Dynamic
Degree
Image
Quality
CogVideoX-2B 96.78 96.63 98.89 97.73 60.82 59.86 61.68
w/o sft loss 96.56 96.83 98.73 97.81 61.25 60.94 61.33
w/ sft loss 96.77 97.01 99.01 97.84 61.55 64.33 62
Table 5: Performance comparison with and without real data training.To verify which model training is more stable, we compared experiments with and without the inclusion of real data. Metrics are evaluated using VBench.
Quantitative results are evaluated using the VBench metric, which comprehensively measures video generation quality in terms of temporal coherence, semantic alignment, and overall plausibility. Our experiments demonstrate that the optimized $2B$ model achieves significant improvements across all evaluated metrics compared to the baseline $5B$ model. Specifically, the optimized model exhibits substantial gains in temporal coherence, effectively enhancing the smoothness of transitions between frames and minimizing abrupt changes. Moreover, the model shows marked improvements in semantic alignment, demonstrating a stronger ability to adhere to the intended meaning of input prompts and better capturing their underlying context. These advancements highlight the efficacy of the proposed post-training strategy in addressing key limitations of the baseline model. The overall VBench score further confirms that the optimized $2B$ model not only achieves superior video generation quality but also maintains computational efficiency, underscoring the practical advantages of our approach in optimizing smaller-scale models to outperform larger counterparts.
5.2.2 Qualitative Results
To further analyze the benefits of our approach, we provide qualitative comparisons between the baseline and optimized models. Visual results demonstrate that the optimized $2B$ model produces videos with higher semantic fidelity, accurately capturing the intent of input prompts and reflecting their underlying context. Additionally, the optimized model ensures greater temporal consistency, effectively minimizing artifacts and sudden transitions between frames, which are often observed in the baseline outputs. Furthermore, the generated videos exhibit enhanced realism, with reduced incoherent structures and improved plausibility, making the outputs more visually appealing and aligned with human perception. These qualitative improvements highlight the effectiveness of our optimization process in addressing key challenges related to semantic and temporal consistency, significantly elevating the overall quality of generated videos.
5.3 Ablation Studies
We conduct a series of ablation studies to validate the contributions of key components in our framework. Each experiment evaluates performance using the VBench metric to ensure consistency in evaluation.
Model Scale Pairwise Pointwise
Accuracy Accuracy
13B 83.72 95.33
40B 86.14 97.57
Table 6: Accuracy of Critic Models on Validation Sets Across Different Scales. The critic models used in this study are finetuned from the ViLA[25] multimodal large language model. This table evaluates their performance at different scales (13B and 40B) in terms of accuracy on pairwise and pointwise validation datasets. The results demonstrate that larger model scales yield higher accuracy for both pairwise ranking and pointwise scoring, indicating improved alignment with human preferences as the model size increases.
5.3.1 Impact of Critic Model Scale on Accuracy
To evaluate the impact of model scale on the performance of the critic model, we conducted experiments with two different scales of ViLA-based critic models: 13B and 40B parameters. The critic models were fine-tuned on a combination of pairwise and pointwise annotated datasets, designed to align with human preferences effectively. Table 6 presents the accuracy of these models on validation datasets for both pairwise ranking and pointwise scoring tasks. The results indicate that increasing the model scale from 13B to 40B leads to significant improvements in both pairwise and pointwise accuracy. Specifically, the 40B model achieves 86.14% accuracy in pairwise ranking and 97.57% in pointwise scoring, outperforming the 13B model by 2.42% and 2.24%, respectively. This demonstrates that larger-scale critic models are better at capturing and generalizing human preferences, likely due to their increased capacity to represent complex relationships in multimodal data. The ablation study highlights the importance of model scale in improving the critic model’s performance. While smaller models can achieve reasonable results, larger models provide a notable boost in accuracy, which is crucial for downstream reinforcement learning tasks. These findings suggest that scaling up the critic model can be a straightforward yet effective way to enhance its alignment capabilities and improve the overall quality of video generation.
5.3.2 Incorporating Real Video Data
To explore the impact of real video data during post-training, we integrated real samples into the training process and applied the original diffusion model loss with reinforcement learning objectives. This modification greatly improves generation quality by enhancing temporal coherence, reducing flickering, and strengthening frame consistency. Furthermore, the inclusion of real-world examples boosts semantic alignment, guiding the model to generate more realistic motion dynamics and contextual details.
As shown in Table 5, the introduction of real video data (indicated by the inclusion of sft loss) leads to measurable improvements across multiple VBench metrics, including temporal flicker, motion smoothness, and aesthetic quality. Notably, the optimized model achieves higher scores in semantic consistency and dynamic degree, indicating a significant enhancement in the realism and coherence of generated videos. These results underscore the effectiveness of integrating real-world data with reinforcement learning to refine video generation outputs, enabling the model to produce videos with superior visual and temporal fidelity.
5.3.3 Comparison of different RL Methods
To evaluate the performance of the proposed online reinforcement learning framework, we conducted experiments using two different feedback alignment methods: Knowledge Transfer Optimization (KTO) and Direct Policy Optimization (DPO). The results, presented in Table 1, demonstrate that our framework consistently improves video generation quality across multiple metrics, regardless of the specific reinforcement learning method employed.
Both KTO and DPO lead to significant enhancements over the baseline CogVideoX-2B model, showing improvements in temporal flicker reduction, motion smoothness, and dynamic degree. These results validate the generality and robustness of the proposed online reinforcement learning framework, which effectively integrates human feedback to optimize video generation outputs. For example, the inclusion of either KTO or DPO raises the Aesthetic Quality and Dynamic Degree metrics compared to the baseline model, underscoring the benefits of aligning the generated outputs with human preferences.
Between the two methods, KTO achieves consistently better performance than DPO across most metrics. Notably, KTO yields a higher Dynamic Degree score (68.83 vs. 68.76) and improved Motion Smoothness (97.93 vs. 98.03), suggesting that KTO is more effective at producing realistic and temporally coherent videos. Additionally, KTO demonstrates greater stability during training, as evidenced by its higher overall scores across Subject Consistency and Background Consistency, which measure adherence to semantic and contextual details. These findings highlight the advantages of the KTO method in leveraging multi-step optimization, enabling the generation of more plausible and visually consistent video outputs compared to DPO.
In summary, the experimental results confirm that the proposed online reinforcement learning framework is effective under different optimization methods. Among these, KTO exhibits superior performance and stability, making it a preferred choice for improving video generation quality through human feedback alignment.
5.3.4 Effect of Multi-Round Training
The impact of multi-round reinforcement learning is assessed by training models with 1, 2, and 3 rounds of iterative optimization, where a single round represents the traditional offline post-training approach. Results in Table 2 show that performance improves with each additional round, as multi-round training better aligns the model with human feedback and refines generation quality. These findings confirm that iterative reinforcement learning effectively enhances both temporal and semantic consistency.
6 Conclusion
This paper presents Iterative Preference Optimization (IPO), a novel post-training framework for text-to-video (T2V) generation. IPO iteratively enhances models using preference datasets, criticism models, and iterative learning, improving subject consistency, motion smoothness, and aesthetic quality. As the first to apply iterative preference optimization to T2V, IPO leverages human feedback to refine video quality, making it more aligned with user expectations. Experiments show that IPO enables a 2B-parameter model to outperform a 5B-parameter model on VBench.